OpenStudy (anonymous):

Question 1: Solve -3 (6x - 4) + 11x = -9

OpenStudy (sleepyjess):

Do you know how to distribute the -3(6x-4) ?

OpenStudy (anonymous):

ya

OpenStudy (sleepyjess):

Okay, so what does the formula become when you do that as your first step?

OpenStudy (anonymous):

-3 (6x - 4) + 11x = -9 -- using the distributive properties, you multiply -3 into the brackets. -18x +12 +11x = -9 -- collect like terms -7x +12 = 9 -- move the positive 12 to the other side -7x = -21 -- divide both sides by -7 to isolate for x ---- ---- -7 -7 x= 3 :)

OpenStudy (anonymous):

is x=3 the answer given to this question? Does my process make sense to you?

OpenStudy (anonymous):

ya @lovethecreation